OpenStudy (anonymous):

Determine whether the point (1, 1) is a solution to the system of equations. Equations: g(x) = 3x + 2 f(x) = (x – 1) + 1…the parenthesis are supposed to represent absolute value brackets. I have plugged in 1 for x in both equations and found that f(1) = 0 and g(1) = 5. Since they don’t match there will be no solution to the system of equations…right? And how do I plug in y (1) into the equations?

OpenStudy (denonakavro):

1,1 is not a solution to the system because 3(1)+2 does not equal 1

OpenStudy (anonymous):

Yes but isn't that just plugging it in for x? How about y?

OpenStudy (denonakavro):

In (1,1) x is 1 and y is 1 right ? So in g(x)= 3x+2, the x will be 1 and the g(x) will be y which is also equal to 1. G(x) basically means y

OpenStudy (anonymous):

So if I plug in 1 for x in the equation, I am solving for y? If I plug in 1 for g(x) or y then I am solving for x?

OpenStudy (denonakavro):

In algebra, when it is f(x) or any other letter g(x), h(x) and etc... that means the y . So to find the answer you have to plug in 1 as y and 1as x 1=3(1)+2 1=5, this is wrong Now lets check the other one 1=(1 – 1) + 1 1=1 this is correct now that the first one is wrong 1,1 is not a solution to the system

OpenStudy (denonakavro):

Hey can you fan me ? Thanks

OpenStudy (anonymous):

Thanks for everything!